# Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness

### v1

@@ 4.1 Experiment Setting
We evaluate on both real-world and synthetic datasets. For vision tasks, we use Waterbirds (sagawadistributionally), CelebA (liu2015faceattributes), MetaShift (liang2022metashift), and ColorMNIST (arjovsky2019invariant). For language tasks, we use CivilComments (borkan2019nuanced) and MultiNLI (williams2017broad). These datasets are widely adopted benchmarks for evaluating robustness to spurious correlations. Detailed descriptions and statistics are in Appendix B.1.1.
@@
Following prior work (gulrajani2020search; izmailov2022feature), we use pretrained ResNet-50 (he2016deep) for image datasets and pretrained BERT (idrissi2022simple) for text datasets. We use SGD with momentum (polyak1964some) for images and AdamW (loshchilov2017decoupled) for text, following standard protocols (yang2023change). Training steps are 5,000 for Waterbirds, MetaShift, and ColorMNIST, and 30,000 for CelebA, CivilComments, and MultiNLI.
@@
We evaluate our method against various bias mitigation and imbalanced learning algorithms. Our comparison includes ERM; subpopulation robustness methods such as GroupDRO (sagawadistributionally), CVaRDRO (duchi2021learning), LfF (nam2020learning), JTT (pmlr-v139-liu21f), LISA (yao2022improving), DFR (kirichenko2022last), PDE (deng2023robust), and ElRep (wen2025elastic); data augmentation approaches including Mixup (zhang2018mixup); domain adaptation methods such as IRM (arjovsky2019invariant) and MMD (li2018domain); and class imbalance techniques including ReSample (japkowicz2000class), ReWeight (japkowicz2000class), SqrtReWeight, Focal Loss (lin2017focal), CB Loss (cui2019class), LDAM (cao2019learning), and Balanced Softmax (ren2020balanced). For ColorMNIST, we evaluate only methods that achieve worst-group accuracy above 70%. Detailed experimental settings are provided in Appendix B.1.2.
@@
ColorMNIST
$\rho=80\%$ $\rho=90\%$ $\rho=95\%$
Algorithm Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc
ERM 38.2 $\pm$2.4 30.9 $\pm$2.7 31.9 $\pm$7.4 7.8 $\pm$3.4 41.1 $\pm$13.0 10.1 $\pm$4.5
GroupDRO 73.5 $\pm$0.3 73.1 $\pm$0.1 73.5 $\pm$0.0 72.7 $\pm$0.3 72.3 $\pm$ 0.1 70.7 $\pm$0.2
LISA 73.8 $\pm$0.1 73.2 $\pm$0.1 73.3 $\pm$0.1 72.9 $\pm$0.2 72.6 $\pm$0.4 71.4 $\pm$0.6
ReSample 73.7 $\pm$0.1 72.2 $\pm$0.3 72.6 $\pm$0.5 70.8 $\pm$1.6 71.8 $\pm$0.1 70.7 $\pm$0.5
PDE 76.6 $\pm$0.8 72.9 $\pm$0.3 72.6 $\pm$0.2 70.2 $\pm$0.7 73.4 $\pm$0.4 70.0 $\pm$0.3
SCER 74.1 $\pm$0.1 73.6 $\pm$0.2 73.6 $\pm$0.1 73.0 $\pm$0.1 73.5 $\pm$ 0.3 72.8 $\pm$ 0.3
@@ 5.1 Quantitative Analysis
Table 1 shows that SCER outperforms baseline methods across four benchmarks: Waterbirds, CelebA, MetaShift, and ColorMNIST, each presenting different spurious correlation challenges. SCER achieves the highest worst-group accuracy in four datasets, with 91.2% on Waterbirds, 91.4% on CelebA, 86.7% on MetaShift, and 73.6% on ColorMNIST at $\rho=80\%$. These results demonstrate that embedding-level disentanglement of spurious and core features effectively mitigates bias.
@@
Table 2 reports performance on ColorMNIST under varying levels of spurious correlation $\rho$. As $\rho$ increases, the dataset becomes increasingly biased, making classification more challenging and highlighting the detrimental impact of spurious correlations on standard models. SCER consistently achieves the highest worst-group accuracy across all values of $\rho$, attaining 73.6% at $\rho=80\%$, 73.0% at $\rho=90\%$, and 72.8% at $\rho=95\%$, outperforming all baselines.
@@
In addition to the standard evaluation settings in Table 2, we examine a more challenging scenario in which a minority group is completely absent during training, creating a label color distribution of 50-0-10-40, as shown in Table 8. This extreme setting tests algorithms’ ability to generalize under harsh spurious correlations when facing complete group absence.
As shown in Table 5.1, SCER achieves 59.6% worst-group accuracy in this extreme setting, significantly outperforming all existing methods. This reveals limitations of existing approaches under extreme spurious correlations. In this extreme setting where an entire subpopulation is omitted from training, GroupDRO (sagawadistributionally) suffers from instability with extreme minority groups, poor extrapolation to unseen
subpopulations, and regularization-capacity trade-offs that limit its effectiveness. class imbalance methods (japkowicz2000class; lin2017focal; cui2019class) and LISA (yao2022improving) fail as they rely on reweighting available samples or interpolating between seen domains, making them unable to handle completely missing groups. Similarly, PDE (deng2023robust), despite initially learning from balanced data, cannot generalize to entirely unseen group combinations. These results demonstrate that under complete group absence, which induces extreme spurious correlations and corrupts the representation space, embedding-level interventions are essential. SCER’s superior performance is attributed to its direct regularization of the representation space, effectively mitigating spurious correlations and enabling robust generalization to entirely unseen subpopulations.
Table 4: Performance comparison without explicit bias labels on ColorMNIST.
@@
A key advantage of SCER is its modular design, enabling seamless integration with existing frameworks without explicit bias labels. We integrate SCER with the Environment Inference for Invariant Learning (EIIL) framework (creager2021environment), replacing EIIL’s second-stage IRM objective with SCER. Table 5.1 shows experimental results integrating EIIL with SCER where no environment information is provided, simulating realistic scenarios without environment labels. We adopt the data setup from (creager2021environment) with two environments of different spurious correlation strengths, detailed in Table 8.
The environments inferred by EIIL achieve over 95% agreement with actual spurious groups in ColorMNIST, providing reliable pseudo-labels for SCER training. As shown in Table 5.1, SCER maintains its effectiveness even with inferred environments, achieving the highest test accuracy of 72.6%. Notably, while GroupDRO suffers significant performance degradation under environment misalignment, SCER remains robust due to its embedding-level regularization approach.
This demonstrates SCER’s practical applicability to real-world scenarios where explicit bias annotations are unavailable.
@@
Table 5.1 presents SCER’s performance on text-based datasets. SCER consistently achieves the best results on both CivilComments for multi-domain evaluation and MultiNLI for multi-class evaluation. On CivilComments, SCER attains the highest worst-group accuracy of 74.0%, effectively mitigating minority subpopulation performance degradation. On MultiNLI, SCER achieves the highest worst-group accuracy of 76.8%, demonstrating balanced learning across different class distributions. These results confirm that SCER’s embedding regularization approach generalizes effectively beyond vision to text-based scenarios. SCER demonstrates robust performance across multi-domain and multi-class environments, highlighting its versatility in handling subpopulation shifts across modalities.
Table 5: Performance comparison on text datasets. from deng2023robust for Civilcomments. †† from wen2025elastic for Civilcomments. SCER achieves the best worst-group accuracy across multi-class and domain settings.
@@
ReWeight 82.5 $\pm$0.0 72.5 $\pm$0.0 81.0 $\pm$0.2 68.8 $\pm$0.4
SqrtReWeight 83.3 $\pm$0.5 71.7 $\pm$0.4 80.7 $\pm$0.3 69.5 $\pm$0.7
CBLoss 82.9 $\pm$0.1 73.3 $\pm$0.2 80.6 $\pm$0.1 72.2 $\pm$0.3
Focal 85.5 $\pm$0.2 62.0 $\pm$1.0 80.7 $\pm$0.2 69.4 $\pm$0.7
LDAM 81.9 $\pm$2.2 37.4 $\pm$8.1 80.7 $\pm$0.3 69.6 $\pm$1.6
BSoftmax 83.8 $\pm$0.0 71.2 $\pm$0.4 80.9 $\pm$0.1 66.9 $\pm$0.4
DFR 83.3 $\pm$0.0 69.6 $\pm$0.2 81.7 $\pm$0.0 68.5 $\pm$0.2
@@ 5.2 Qualitative Analysis
We conduct sensitivity analysis on ColorMNIST with $\rho=95\%$ to evaluate our proposed loss components. As shown in Table 6, we systematically vary $\lambda_{\text{core}}$ and $\lambda_{\text{spur}}$ according to Equation 2. Results show each component independently improves worst-group accuracy, while their joint optimization achieves the highest robustness, confirming that $\mathcal{L}_{\text{core}}$ and $\mathcal{L}_{\text{spur}}$ produce complementary effects, validating our theoretical motivation. Additional analysis on CelebA is in Appendix C.
@@
We investigate our proposed $\Sigma$-norm, also known as the Mahalanobis distance, by replacing the standard Euclidean norm in Equation 1. As shown in Table 5.2, the $\Sigma$-norm outperforms the Euclidean baseline. This improvement comes from the natural benefits of Mahalanobis distance, which captures the structure of covariance of the features and ensures the invariance of the scale in high-dimensional spaces (ghorbani2019mahalanobis; mahalanobis2018generalized).
@@
Figure 3 demonstrates the relationship between worst-group accuracy and spurious/core metrics, revealing insights into group-invariant learning. Spurious Loss exhibits negative correlations with Worst Acc, indicating that stronger classifier dependence on domain-specific characteristics degrades minority group performance, while Core Loss shows positive correlations, demonstrating that alignment with genuine label-discriminative features enhances cross-group generalization. Specifically, Weight-Spurious Alignment $\text{cor}(\beta^{*},\Delta_{\text{spur}})$ measures how much classifier weights align with intra-class domain variations, quantifying spurious feature dependence, while Weight-Core Alignment $\text{cor}(\beta^{*},\Delta_{\text{core}})$ captures how well the classifier aligns with domain-consistent discriminative directions. Spurious Magnitude $\|\Delta_{\text{spur}}\|_{\Sigma}$ and Core Magnitude $\|\Delta_{\text{core}}\|_{\Sigma}$ quantify inter-domain distributional differences and inter-label separability, respectively. These results establish that balancing core signal capture and spurious correlation suppression is fundamental to group invariance.
@@
We compare ERM and SCER using Grad-CAM in Figure 4. ERM focuses on background regions, indicating spurious cue reliance. SCER consistently focuses on the bird itself, demonstrating that embedding regularization redirects attention to the true object and produces robust representations.
@@ 6 Conclusion
We propose Spurious Correlation-Aware Embedding Regularization (SCER), which addresses spurious correlations through embedding-level regularization that decomposes worst-group error into spurious and core components. SCER penalizes spurious feature alignment while promoting core feature consistency across domains, thereby mitigating spurious correlation overfitting. Experiments demonstrate superior worst-group accuracy, confirming SCER’s robust performance.

### v2

@@ 4.1 Experiment Setting
We evaluate on both real-world and synthetic datasets. For vision tasks, we use Waterbirds (Sagawa et al., ), CelebA (Liu et al., 2015), MetaShift (Liang and Zou, 2022), and ColorMNIST (Arjovsky et al., 2019). For language tasks, we use CivilComments (Borkan et al., 2019) and MultiNLI (Williams et al., 2018). These datasets are widely adopted benchmarks for evaluating robustness to spurious correlations. Detailed descriptions and statistics are in Appendix B.1.1.
@@
Following prior work (Gulrajani and Lopez-Paz, ; Izmailov et al., 2022), we use pretrained ResNet-50 (He et al., 2016) for image datasets and pretrained BERT (Idrissi et al., 2022) for text datasets. We use SGD with momentum (Polyak, 1964) for images and AdamW (Loshchilov and Hutter, 2017) for text, following standard protocols (Yang et al., 2023b). Training steps are 5,000 for Waterbirds, MetaShift, and ColorMNIST, and 30,000 for CelebA, CivilComments, and MultiNLI.
@@
We evaluate our method against various bias mitigation and imbalanced learning algorithms. Our comparison includes ERM; subpopulation robustness methods such as GroupDRO (Sagawa et al., ), LISA (Yao et al., 2022), PDE (Deng et al., 2023), and ElRep (Wen et al., 2025); and class imbalance techniques such as ReSample (Japkowicz, 2000). For ColorMNIST, we evaluate only methods that achieve worst-group accuracy above 70%. We conduct experiments on a broader set of baselines, with detailed descriptions and complete results provided in Appendix B.1.2 and Tables 11, 12, C.1, and C.1. We also conduct a focused comparison against GroupDRO-ES (Izmailov et al., 2022), the most competitive GroupDRO variant with early stopping, as presented in Table 16.
@@ 5.1 Quantitative Analysis
Table 1 shows that SCER outperforms baseline methods across four benchmarks: Waterbirds, CelebA, MetaShift, and ColorMNIST, each presenting different spurious correlation challenges. SCER achieves the highest worst-group accuracy in four datasets, with 91.2% on Waterbirds, 91.4% on CelebA, 86.7% on MetaShift, and 73.6% on ColorMNIST at $\rho=80\%$. These results demonstrate that embedding-level disentanglement effectively mitigates spurious bias.
@@
Table 2 reports performance on ColorMNIST under varying levels of spurious correlation $\rho$. As $\rho$ increases, the dataset becomes increasingly biased, making classification more challenging and highlighting the detrimental impact of spurious correlations on standard models. SCER consistently achieves the highest worst-group accuracy across all values of $\rho$, attaining 73.6% at $\rho=80\%$, 73.0% at $\rho=90\%$, 72.8% at $\rho=95\%$, and 56.0% at $\rho=99\%$, outperforming all baselines.
ColorMNIST
$\rho=80\%$ $\rho=90\%$ $\rho=95\%$ $\rho=99\%$
Algorithm Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc Avg Acc Worst Acc
ERM 38.2 $\pm$2.4 30.9 $\pm$2.7 31.9 $\pm$7.4 7.8 $\pm$3.4 41.1 $\pm$13.0 10.1 $\pm$4.5 50.7 $\pm$8.2 8.5 $\pm$5.0
GroupDRO 73.5 $\pm$0.3 73.1 $\pm$0.1 73.5 $\pm$0.0 72.7 $\pm$0.3 72.3 $\pm$ 0.1 70.7 $\pm$0.2 50.3 $\pm$3.9 38.7 $\pm$0.8
LISA 73.8 $\pm$0.1 73.2 $\pm$0.1 73.3 $\pm$0.1 72.9 $\pm$0.2 72.6 $\pm$0.4 71.4 $\pm$0.6 53.1 $\pm$7.2 10.2 $\pm$8.1
ReSample 73.7 $\pm$0.1 72.2 $\pm$0.3 72.6 $\pm$0.5 70.8 $\pm$1.6 71.8 $\pm$0.1 70.7 $\pm$0.5 57.1 $\pm$3.0 45.4 $\pm$10.0
PDE 76.6 $\pm$0.8 72.9 $\pm$0.3 72.6 $\pm$0.2 70.2 $\pm$0.7 73.4 $\pm$0.4 70.0 $\pm$0.3 56.3 $\pm$4.3 47.5 $\pm$8.9
SCER 74.1 $\pm$0.1 73.6 $\pm$0.2 73.6 $\pm$0.1 73.0 $\pm$0.1 73.5 $\pm$0.3 72.8 $\pm$0.3 61.0 $\pm$2.0 56.0 $\pm$2.2
@@
In addition to the standard evaluation settings in Table 2, we examine a more challenging scenario in which a minority group is completely absent during training, creating a label color distribution of 50-0-10-40, as shown in Table 8. This extreme setting tests algorithms’ ability to generalize under harsh spurious correlations when facing complete group absence. As shown in Table 5.1, SCER achieves 59.6% worst-group accuracy in this extreme setting, significantly outperforming all existing methods. This result reveals limitations of existing approaches under extreme spurious correlations.
In this extreme setting where an entire subpopulation is omitted from training, GroupDRO (Sagawa et al., ) suffers from instability with extreme minority groups, poor extrapolation to unseen subpopulations, and regularization-capacity trade-offs that limit its effectiveness. Class imbalance methods (Japkowicz, 2000) and LISA (Yao et al., 2022) fail as they rely on reweighting available samples or interpolating between seen domains, making them unable to handle completely missing groups. Similarly, PDE (Deng et al., 2023), despite initially learning from balanced data, cannot generalize to entirely unseen group combinations. These results show embedding-level interventions are essential when complete groups are absent, as this causes extreme spurious correlations. SCER’s superior performance stems from directly regularizing the representation space, mitigating spurious correlations and enabling robust generalization to unseen subpopulations.
Table 4: Performance comparison without explicit bias labels.
@@
A key advantage of SCER is its modular design, which enables seamless integration with existing frameworks without requiring explicit bias labels. We integrate SCER with the Environment Inference for Invariant Learning (EIIL) framework (Creager et al., 2021), replacing EIIL’s second-stage IRM objective with SCER. Table 5.1 shows experimental results integrating EIIL with SCER where no environment information is provided, simulating realistic scenarios without environment labels.
We adopt the data setup from (Creager et al., 2021), with two environments differing in the strength of spurious correlations, as detailed in Table 8. The environments inferred by EIIL achieve over 95% agreement with actual spurious groups in ColorMNIST, providing reliable pseudo-labels for SCER training. As shown in Table 5.1, SCER maintains its effectiveness even with inferred environments, achieving the highest test accuracy of 72.6%. Notably, while GroupDRO suffers significant performance degradation under environmental misalignment, SCER remains robust due to its embedding-level regularization. This demonstrates SCER’s practical applicability and adaptability to real-world scenarios where explicit bias annotations are unavailable. Even in scenarios where environment labels are increasingly noisy, SCER continues to demonstrate robust performance, as detailed in Table C.1.
@@
Table 5.1 presents SCER’s performance on text-based datasets. SCER consistently achieves the best results on both CivilComments for multi-domain evaluation and MultiNLI for multi-class evaluation.
On CivilComments, SCER achieves the highest worst-group accuracy of 74.0%, effectively mitigating performance degradation in minority subpopulations. On MultiNLI, SCER achieves the highest worst-group accuracy of 76.8%, demonstrating balanced learning across different class distributions. These results confirm that SCER’s embedding regularization approach generalizes effectively beyond vision to text-based scenarios.
Table 5: Performance comparison on text datasets. from Deng et al. (2023) for Civilcomments. †† from Wen et al. (2025) for Civilcomments. SCER achieves the best worst-group accuracy across multi-class and domain settings.
@@
SCER demonstrates robust performance across multi-domain and multi-class environments, highlighting its versatility in handling subpopulation shifts across modalities.
Our quantitative analysis shows that SCER consistently surpasses existing methods across diverse image and text settings, maintaining strong worst-group performance even under severe spurious correlations, missing groups, and inferred environments. This demonstrates that embedding-level regularization effectively reduces spurious bias and supports robust generalization across domains.
@@ 5.2 Qualitative Analysis
We conduct sensitivity analysis on ColorMNIST with $\rho=95\%$ to evaluate our proposed loss components. As shown in Table 6, we systematically vary $\lambda_{\text{core}}$ and $\lambda_{\text{spur}}$ according to Equation 2. Results show that each regularization component independently improves worst-group accuracy over the baseline, demonstrating that both $\mathcal{L}_{\text{core}}$ and $\mathcal{L}_{\text{spur}}$ contribute to robustness against spurious correlations. Notably, their joint optimization achieves the highest worst-group performance, confirming that the two loss terms produce complementary effects. This validates our theoretical insight that achieving optimal worst-group generalization requires simultaneously maximizing alignment with core features while minimizing reliance on spurious features. To further assess parameter sensitivity in a large-scale setting, we provide additional analysis on CelebA in Table 18.
Figure 3: t-SNE visualization on the Waterbirds dataset. ERM clusters samples primarily by background rather than label. GroupDRO partially mitigates this but still exhibits background-based separation within each label. SCER produces label-aligned, background-invariant embeddings, effectively suppressing spurious correlations.
Embedding Space Analysis
Figure 3 shows t-SNE projections of embeddings learned by the three models on the Waterbirds dataset. ERM exhibits strong domain-based clustering, with samples separating predominantly by background rather than label. Samples from different backgrounds form distinct clusters, indicating heavy reliance on spurious features rather than core class characteristics. GroupDRO shows reduced but still visible clustering by background within the same label, demonstrating incomplete suppression of spurious correlations.
In contrast, SCER produces label-aligned embeddings where samples cluster primarily by label regardless of background. Waterbird samples from different backgrounds intermix to form a single unified cluster, while Landbird samples from different backgrounds aggregate similarly, maintaining clear separation between the two classes. This domain-invariant structure provides qualitative evidence that SCER effectively suppresses spurious correlations via embedding-level regularization, resulting in representations that generalize robustly across distribution shifts.
@@
We investigate our proposed $\Sigma$-norm by replacing the standard Euclidean norm in Equation 1. As shown in Table 5.2, the $\Sigma$-norm outperforms the Euclidean baseline. This improvement comes from the natural benefits of the distance, which captures the structure of covariance of the features and ensures the invariance of the scale in high-dimensional spaces (Ghorbani, 2019; Mahalanobis, 2018).
@@
Figure 4 demonstrates the relationship between worst-group accuracy and spurious/core metrics, revealing insights into group-invariant learning. Spurious Loss exhibits negative correlations with Worst Acc, indicating that a stronger dependence of the classifier on domain-specific characteristics degrades the performance of the minority group, while Core Loss shows positive correlations, demonstrating that alignment with genuine label discriminatory features enhances the generalization of the group.
Specifically, Weight-Spurious Alignment $\text{cor}(\beta^{*},\Delta_{\text{spur}})$ measures how much classifier weights align with intra-class domain variations, quantifying spurious feature dependence. In contrast, Weight-Core Alignment $\text{cor}(\beta^{*},\Delta_{\text{core}})$ captures how well the classifier aligns with domain-consistent discriminative directions. Spurious Magnitude $\|\Delta_{\text{spur}}\|_{\Sigma}$ and Core Magnitude $\|\Delta_{\text{core}}\|_{\Sigma}$ quantify inter-domain distributional differences and inter-label separability, respectively. The consistent patterns observed across these metrics align with our theoretical decomposition, confirming that the regularization terms directly address the properties essential for worst-group robustness. These results establish that balancing core signal capture and spurious correlation suppression is fundamental to group invariance.
Figure 5: Grad-CAM visualization comparison on the Waterbirds. ERM heavily relies on background cues, and GroupDRO only partially captures the object, whereas SCER consistently attends to bird-specific features. This demonstrates SCER’s ability to suppress spurious background correlations while focusing on label-relevant structures.
@@
Figure 5 presents a comparative analysis of Grad-CAM (Selvaraju et al., 2017) visualizations for three methods applied to bird classification models. ERM exhibits dispersed activation patterns across the entire image, allocating considerable attention to background regions. GroupDRO shows concentrated activation on some regions of the bird, but does not fully utilize all parts that can be considered core features for classification. In contrast, SCER concentrates strong activation across structural features such as the bird’s body, head, wings, and tail, while effectively suppressing dependence on background or spurious elements. These results visually demonstrate that our mean embedding-based penalty approach comprehensively captures diverse discriminative features critical for classification while effectively suppressing spurious factors. Additional examples are provided in Figure 6.
Our qualitative analysis demonstrates that SCER effectively suppresses spurious features and learns representations focused on core signals, as evidenced by embedding structures, attention visualizations, and interpretability analyses that support robust group-invariant behavior.
@@ 6 Conclusion
We propose Spurious Correlation-Aware Embedding Regularization (SCER), which addresses spurious correlations by decomposing worst-group error into spurious and core components at the embedding level. By suppressing alignment with spurious feature directions while promoting consistency with core feature directions, the model is prevented from overfitting to domain-specific artifacts. Through extensive experiments across image and text domains, we confirm that SCER consistently achieves superior worst-group accuracy, demonstrating that explicit embedding-level regularization addresses subgroup bias more directly than sample reweighting or feature alignment approaches. Furthermore, SCER’s modular design enables integration with environment-inference methods, making it practically applicable to real-world scenarios without requiring explicit bias labels. This work presents a promising direction for robust learning under distribution shift and is expected to contribute to practical applications where fairness is critical.
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216).